OpenStudy (anonymous):

what is the value of f(5) for the function f(x)=4x-3

OpenStudy (smored):

replace x with 5 4(5)-3=17

OpenStudy (anonymous):

f(x) = 4x - 3 f(5) = 4(5) - 3 f(5) = 20 - 3 f(5) = 17 Replace 5 for every x in the function

OpenStudy (anonymous):

im sorry but im still lost

OpenStudy (anonymous):

just replace x for 5

OpenStudy (smored):

and solve it

OpenStudy (anonymous):

f(x) --> f(5) so 5 took the x's place therefore replace every x for 5

OpenStudy (anonymous):

Get it hun?

OpenStudy (anonymous):

sorta im a little on the slow side sadly

OpenStudy (anonymous):

Hm, k. So it says that f(x) is f(5) So, in f(x) = 4x - 3 I replaced 5 for x, so in f(x) it becomes f(5) and because 4 is being multiplied by x then I mulitply 4 * 5

OpenStudy (mathmale):

Barbie, Sorry this is confusing. What you have here, f(x)=4x-3, is called a "function." It tells us what value y=f(x) takes on at the input value x. "f(5)" requires that we substitute 5 for x in the function. Thus, f(x)=4X-3 becomes f(5)=4(5)-3 when we substitute 5 for x. Thus, the function value at x=5 is 4(5)-3, which can be reduced to 20-3, or 17.

OpenStudy (mathmale):

Ello's approach is similar and is correct, except that the very last sentence should include, "then I multiply 4*5 and subtract 3 from the result."

OpenStudy (mathmale):

Which of the answer options is the correct one?